Greg started with a certain number of quarters. He then decided on a number of quarters he would save each day. He added the quarters he saved to the amount with which he started.
At the end of day 2, Greg had a total of 26 quarters saved.
At the end of day 5, he had a total of 35 quarters saved.
At the end of day 8, he had a total of 44 quarters saved.
Hint: Create an x/y or t-chart and fill in the numbers with x-value representing the # of days and the y-value representing "total coins."
A. How many quarters does Greg start with (day zero)? Show or explain your work.
B. Write an equation in slope-intercept form to model the total quarters Greg has saved after x days.
C. Using the rate at which Greg is saving, explain why he can never have exactly 100 quarters saved by the end of any given day.

Answer by greenestamps(13334)



He is adding 9 quarters every 3 days, or 3 quarters every day.

A. At the end of day 2, he had added 2*3=6 quarters to the starting number. Since he had 26 quarters at the end of day 2, he started with 26-6=20 quarters.

B. 20, plus 3 more for each day: literally, y = 20+3x. In slope-intercept form, y = 3x+20.

C. The starting number of 20 is 2 more than a multiple of 3. Since he adds 3 quarters each day, the total at the end of every day will be 2 more than a multiple of 3. 100 is 1 more than a multiple of 3, so he will never have 100 quarters at the end of any day.

Note that instead of explaining part C in words, you could just show that the equation 3x+20=100 has a non-integer solution.